# A Comprehensive Survey of Deep Learning for Multivariate Time Series Forecasting: A Channel Strategy Perspective

### v2

@@ Abstract
Multivariate Time Series Forecasting (MTSF) plays a crucial role across diverse fields, ranging from economic, energy, to traffic. In recent years, deep learning has demonstrated outstanding performance in MTSF tasks. In MTSF, modeling the correlations among different channels is critical, as leveraging information from other related channels can significantly improve the prediction accuracy of a specific channel. This study systematically reviews the channel modeling strategies for time series and proposes a taxonomy organized into three hierarchical levels: the strategy perspective, the mechanism perspective, and the characteristic perspective. On this basis, we provide a structured analysis of these methods and conduct an in-depth examination of the advantages and limitations of different channel strategies. Finally, we summarize and discuss some future research directions to provide useful research guidance. Moreover, we maintain an up-to-date Github repository [^1] https://github.com/decisionintelligence/CS4TS which includes all the papers discussed in the survey.
@@ 1 Introduction
Multivariate time series forecasting (MTSF) is a fundamental yet challenging task in various domains, including economic, energy, and traffic Qiu et al. (2024); Wu et al. (2024); Yu et al. (2024). The ability to accurately predict future values of multiple interdependent channels (a.k.a., variables) over time is crucial for making informed decisions, optimizing resource allocation, and improving operational efficiency.
@@
While there are several surveys on MTSF Wen et al. (2023); Wang et al. (2024d), they often lack a comprehensive discussion on the role of channel strategies in multivariate settings. This study aims to bridge this gap by summarizing the main developments of channel strategies of MTSF. We first briefly introduce the MTSF task and propose a new taxonomy organized into three hierarchical levels. Starting with the strategy perspective, we systematically introduce the definitions of three channel strategies and their representative methods, providing researchers with a foundational framework to understand these strategies. Next, from the mechanism perspective, we further explore how each method implements the specific mechanisms of channel strategies, categorizing and summarizing them to help researchers gain a deeper understanding of the underlying implementation of different channel strategies. Finally, from the characteristic perspective, we focus on the different characteristics considered by channel strategies when modeling the correlations among channels. Following this, we provide a detailed comparison of the advantages and limitations of the three channel strategies, offering valuable insights for researchers. In conclusion, we discuss future research directions that will provide useful guidance for the further development of the field. To the best of our knowledge, this is the first work to comprehensively and systematically review the key developments of deep learning methods for MTSF through the lens of channel strategy. In summary, the main contributions of this survey include:
@@
  • Future research oppotunities: We discuss and highlight future avenues for enhancing MTSF through diverse channel strategies, urging researchers to delve deeper into this area.
@@ 2 Preliminaries
Multivariate Time Series Forecasting: Given a historical multivariate time series $X\in\mathbb{R}^{T\times N}$ of $T$ time points, multivariate time series forecasting aims to predict the next $F$ future time points, i.e., $Y\in\mathbb{R}^{F\times N}$, where $F$ is called the forecasting horizon.
@@
- - - - - - - Chronos Ansari et al. (2024) Foundation ICML 2024 Chronos
- - - - - - - LLM4TS Chang et al. (2023) Foundation NIPS 2023 LLM4TS
@@
GNN-based - - Ada-MSHyper Shang et al. (2024) Specific NIPS 2024 Ada-MSHyper
@@
Asym.: Asymmetry, Lag.: Lagginess, Pol.: Polarity, Gw.: group-wise, Dyn.: Dynamism, and Ms.: Muti-scale. We will discuss them in Section 3.3.
@@ 3.1 Strategy Perspective
Channel Partiality (CP): The CP strategy strikes a balance between CI and CD, allowing each channel to retain a degree of independence while simultaneously interacting with other related channels. This approach emphasizes a hybrid state where channels selectively interact and exhibit partial correlations. Based on whether the number of related channels for each channel is fixed or dynamic, the existing CP methods can be divided into two categories: I) Fixed partial channels: These models fix the number of related channels for each channel, which means the set of associated channels remains contant overtime. For example, in MTGNN Wu et al. (2020), the channel relationships are modeled as a K-regular graph, where each channel interacts with $K$ other channels using the CD strategy to model interdependencies, while the remaining channels interact through the CI strategy. Similarly, in MCformer Han et al. (2024b), each channel interacts with only $K$ other channels, maintaining the CI strategy with the rest to ensure computational efficiency and prevent overfitting. II) Dynamic partial channels: These models allow the number of related channels for each channel to be dynamic, changing over time and providing greater flexibility to adapt to varying scenarios. For instance, DUET Qiu et al. (2025) calculates channel similarity using metric learning in the frequency domain and then sparsifies the result. This creates a mask matrix, which is integrated into the attention mechanism of the fusion module, ensuring that each channel interacts only with relevant channels, reducing interference from noisy ones. Another example, CCM Chen et al. (2024), dynamically clusters channels based on their intrinsic similarities. To effectively capture the underlying time series patterns within these clusters, CCM utilizes a cluster-aware feed forward mechanism, enabling tailored management and processing for each individual cluster.
@@ 3.2 Mechanism Perspective
Transformer-based: In recent years, Transformer has been widely applied to MTSF tasks, leveraging its powerful global modeling capability to effectively capture complex temporal dependencies and channel interactions. As shown in Figure 2, existing channel strategies based on the attention mechanism can be categorized into the following types. I) Naive Attention: These approaches all adopt CD strategy, treating time series segments (patches) or the entire sequence of each channel as individual tokens, and directly applying attention mechanisms to model channel correlations. For instance, CARD Wang et al. (2024b) and iTransformer Liu et al. (2024b) represent the patches and series of each channel as independent tokens, respectively, and explicitly capture channel correlation using attention mechanisms. II) Router Attention: When the number of channels ($N$) is large, the computational complexity of channel attention reaches $O(N^{2})$, resulting in high computational costs. To address this, some methods propose optimization strategies to mitigate the computational complexity caused by CD strategy. For example, Crossformer Zhang and Yan (2022) introduces a Router Mechanism for Naive Attention, which uses a small fixed number of $c$ “routers" ($c\ll N$) to gather information from all channels and redistribute it. This reduces the complexity to $O(2cN)=O(N)$. This mechanism effectively balances the modeling of channel correlation and computational efficiency. III) Frequency Attention: Some CD methods suggest that frequency-domain information is more effective for capturing inter-channel dependencies than time-domain information. For example, FECAM Jiang et al. (2023) transforms the time series data into the frequency domain and then employs Naive Attention to model inter-channel relationships in this domain. IV) Mask Attention: In naive attention, each channel calculates attention scores with all channels, which can be negatively affected by irrelevant channels. To mitigate this, Mask Attention provides an approach to avoid irrelevant noise by constructing CP strategy. For example, DUET Qiu et al. (2025) generates mask matrices for Naive Attention, allowing each channel to focus on those beneficial for downstream prediction tasks, while mitigating the impact of noisy or irrelevant channels. This approach explicitly constrains the attention computation, improving the accuracy of channel correlation modeling.
@@
CNN-based: Convolutional Neural Network (CNN) is deep learning model that utilize convolutional layers to extract local features from data. As illustrated in Figure 3, the explored CNN-based approaches can be broadly categorized as follows: I) Merging: Many models, such as Informer Zhou et al. (2021), Autoformer Wu et al. (2021), and FEDformer Zhou et al. (2022), use 1D convolution with a sliding operation along the temporal dimension in the initial feature extraction layers. These models treat different channels as distinct inputs to the convolution, whose features are then weighted and merged during the convolution process, enabling inter-channel interactions. Although TimesNet Wu et al. (2023a) employs 2D convolutions, it folds the temporal dimension into a 2D format, with variable channels still serving as independent input for weighted merging via convolution. Such models are all under the CD strategy. II) Convolution: Given the slight spatial dependence among channels, ModerTCN Luo and Wang (2024) directly applies convolution operations to facilitate information interaction among channels within local scopes. Within the same convolution window, channels interact with each other in a CD manner through the convolution kernel, while channels that cannot be assigned to the same window remain independent of each other. This results in an efficient method for CP modeling.
GNN-based: By dividing the time series into different windows along time, where each channel within a window is treated as a node and the correlations between channels are considered as edges, multivariate time series can be transformed into graph-based data. The GNN-based methods can be classified into dense and sparse graphs. In dense graphs, each node is typically connected to almost all other nodes, with the edges often representing the strength of correlation or the probability of correlated influence. Methods based on dense graphs, such as GTS Shang et al. (2021) and FourierGNN Yi et al. (2023), generally follow a CD strategy. In contrast, sparse graphs only retain necessary edges, with most nodes remaining independent. For instance, MTGNN Wu et al. (2020) preserves K edges per node, constructing a sparse K-regular graph. Different from this, MTSF-DG Zhao et al. (2023) sparsifies the adjacency matrix by filtering out low-probability edges based on a pre-set threshold. Methods based on sparse graphs belong to the CP strategy.
@@
The sparsity of the constructed graph determines whether the method follows a CD or CP strategy. Additionally, GNN-based models often rely on the type of graph they construct when implementing the CD or CP strategy. As shown in Figure 4, we classify graph types as follows: I) Simple Graph: A simple graph is the most basic graph model, where there is at most one edge between each pair of nodes. A well-defined graph structure is required for effective message passing. Researchers have used channel similarity metrics (MTGNN, MSGNet Cai et al. (2024), CrossGNN Huang et al. (2023a)) and data similarity metrics (GTS, WaveForM Yang et al. (2023)) to learn the correlation graph structure among multivariate channels. They utilize time-domain (MTGNN, MSGNet, CrossGNN, GTS) or frequency-domain (WaveForM) information as node learning features. Graph convolution-based message passing is applied within the simple graph to facilitate the transmission of dependency information among channels. II) Spatio-temporal Graph: Unlike a simple graph, a Spatio-temporal Graph incorporates multiple channels at different time steps into a single graph, further considering the relationships between channels across different time steps. This approach allows GNNs to simultaneously model both temporal and channel dependencies, effectively addressing potential compatibility issues between the temporal module and the GNNs. The main challenge of Spatio-temporal Graph-based methods is to address the efficiency issues in the graph construction and message passing stages. For example, FourierGNN uses fully connected graph construction and employs Fourier domain convolution operators to achieve a time complexity of $O(Nlog(N))$. Similarly, FC-STGNN Wang et al. (2024c) adopts the same graph construction method and employs moving-pooling convolution to achieve the same time complexity. III) Hyper Graph: Hypergraphs are an extension of graphs that allow hyperedges to connect multiple vertices, enabling the modeling of higher-order group interactions. Models based on hypergraphs assume that the interactions among channels are not pairwise but involve group-based interactions among multiple channels. Therefore, hypergraph-based models are inherently suitable for constructing CP strategies. ReMo Wu et al. (2023b) and Ada-MSHyper construct multi-view and multi-scale hypergraphs, respectively, and design message passing mechanisms on these hypergraphs to enable group-wise message propagation. It is noteworthy that they use different MLPs or clustering constraints to promote the expression of heterogeneity among groups. IV) Temporal Graph: In the real world, the correlation of time series data often changes over time, forming dynamic relational graphs. MTSF-DG and TPGNN Liu et al. (2022) use dynamic graphs and polynomial graphs, respectively, to model the variation patterns of these correlations. The CP model MTSF-DG combines historical and future relational graphs, leveraging memory networks and logical symbol learning to capture the impact of historical correlations on future correlations. The CD model TPGNN represents the correlation matrix as a matrix polynomial with time-varying coefficients to learn the evolving patterns of correlations.
@@
Others: In addition to the mechanisms mentioned above, some models have proposed alternative approaches. For example: I) The CD model SOFTS Han et al. (2024a) introduces the STAR module, which utilizes a centralized structure to first aggregate information from all channels using MLPs, and then distribute the aggregated information to each channel. This interaction not only reduces the complexity of inter-channel interactions but also minimizes reliance on individual channel quality. II) The CP model LIFT Zhao and Shen (2024) proposes a novel plugin, adaptable to all MTSF specific models, that efficiently estimates leading indicators and their lead steps at each time step. This approach enables lagging channels to utilize advanced information from a predefined set of leading indicators. III) C-LoRA Nie et al. (2024) introduces a channel-aware low-rank adaptation (C-LoRA) plugin, which is adaptable to all MTSF specific models. It first parameterizes each channel with a low-rank factorized adapter to enable individualized treatment. The specialized channel adaptation is then conditioned on the series information to form an identity-aware embedding. Additionally, cross-channel relational dependencies are captured by integrating a globally shared CD model.
@@ 3.3 Characteristics Perspective
Lagginess: Lagginess refers to the fact that the current state of a certain channel not only depends on the current states of the other channels but may also be influenced by the past states of the other channels. Based on Lagginess characteristic, VCformer Yang et al. (2024) incorporates the joint effects of multi-step delays among channels when calculating the attention matrix. In contrast, FourierGNN Yi et al. (2023) and FC-STGNN Wang et al. (2024c) directly perform message passing between representations across different channels and time steps using spatiotemporal fully connected graphs. LIFT Zhao and Shen (2024), on the other hand, combines prior knowledge with neural network predictions to estimate the lag step.
@@
Group-wise: Group-wise refers to a phenomenon in which correlations among channels exhibit a grouping structure, characterized by strong correlations within the same group, weak correlations among different groups, and varying correlation dependencies across different groups. CCM Chen et al. (2024) and DUET Qiu et al. (2025) use clustering techniques to group channels for interaction, while ReMo Wu et al. (2023b) and Ada-MSHyper Shang et al. (2024) establish intra-group message passing through hyperedges. Furthermore, CCM and ReMo apply different MLPs for feature extraction within different groups, and Ada-MSHyper constrains hyperedges based on the loss function. These varying approaches facilitate the expression of differences among different groups.
Dynamism: The correlation among channels in multivariate time series exhibits different behaviors at different time steps, showing an overall dynamic change. First, methods based on MLP (such as TimeMixer Wang et al. (2024a), TTM Ekambaram et al. (2024)), where the weights remain constant across time steps, fail to capture dynamism. Methods that use Transformer to consider channel correlation typically employ series tokens or patch tokens. Methods based on series tokens, such as iTransformer Liu et al. (2024b) and DUET Qiu et al. (2025), cannot capture dynamism. However, methods based on patch tokens, such as Crossformer Zhang and Yan (2022), assign different attention scores at different time patches, enabling the modeling of dynamism. In GNNs, only approaches where the graph structure changes over time can capture dynamism, such as MSGNet Cai et al. (2024), which computes the graph structure at each timestep. However, the aforementioned methods for modeling dynamism only consider different channel relationships at different time steps. In contrast, MSTF-DG Zhao et al. (2023), TPGNN Liu et al. (2022), and ESG Ye et al. (2022) propose that there is a direct connection between channel relationships across different time steps. For example, MSTF-DG uses previous channel relationships to directly infer the current channel relationships.
Muti-scale: Multi-scale refers to the phenomenon where the correlations among channels exhibit different behaviors at various time scales (such as hours, minutes, or seconds). MSGNet Cai et al. (2024) and Ada-MSHyper Shang et al. (2024) establish different graph structures across scales to describe the variations in correlation at different levels, and they achieve the fusion of correlation information at different scales through varying degrees of interaction. Considering the multi-scale heterogeneity of correlations helps the model better understand the multi-scale features of time series data, thereby generating more accurate predictions.
Table 2: Comparison among different channel strategies.
Dimension CI CD CP
Efficiency High Low Moderate
Robustness High Low Moderate
Generalizability Low Moderate High
Capacity Low High Moderate
Ease of Implementation High Moderate Low

### v3

@@ Abstract
Multivariate Time Series Forecasting (MTSF) plays a crucial role across diverse fields, ranging from economics to energy to traffic. In recent years, deep learning has demonstrated outstanding performance in MTSF tasks. In MTSF, modeling the correlations among different channels is critical, as leveraging information from other related channels can significantly improve the prediction accuracy of a specific channel. This study systematically reviews the channel modeling strategies for time series and proposes a taxonomy organized into three hierarchical levels: the strategy perspective, the mechanism perspective, and the characteristic perspective. On this basis, we provide a structured analysis of these methods and conduct an in-depth examination of the advantages and limitations of different channel strategies. Finally, we summarize and discuss some future research directions to provide useful research guidance. Moreover, we maintain an up-to-date GitHub repository [^1] https://github.com/decisionintelligence/CS4TS which includes all the papers discussed in the survey.
@@ 1 Introduction
Multivariate time series forecasting (MTSF) is a fundamental yet challenging task in various domains, including economic, energy, and traffic Qiu et al. (2024). The ability to accurately predict future values of multiple interdependent channels (a.k.a., variables) over time is crucial for making informed decisions, optimizing resource allocation, and improving operational efficiency.
@@
While there are several surveys on MTSF Wen et al. (2023); Wang et al. (2024d), they often lack a comprehensive discussion on the role of channel strategies in multivariate settings. This study aims to bridge this gap by summarizing the main developments of channel strategies of MTSF. We first briefly introduce the MTSF task and propose a new taxonomy organized into three hierarchical levels. Starting with the strategy perspective, we systematically introduce the definitions of three channel strategies and their representative methods, providing researchers with a foundational framework to understand these strategies. Next, from the mechanism perspective, we further explore how each method implements the specific mechanisms of channel strategies, categorizing and summarizing them to help researchers gain a deeper understanding of the underlying implementation of different channel strategies. Finally, from the characteristic perspective, we focus on the different characteristics considered by channel strategies when modeling the correlations among channels. Following this, we provide a detailed comparison of the advantages and limitations of the three-channel strategies, offering valuable insights for researchers. In conclusion, we discuss future research directions that will provide useful guidance for the further development of the field. To the best of our knowledge, this is the first work to comprehensively and systematically review the key developments of deep learning methods for MTSF through the lens of channel strategy. In summary, the main contributions of this survey include:
@@
  • Future research opportunities : We discuss and highlight future avenues for enhancing MTSF through diverse channel strategies, urging researchers to delve deeper into this area.
@@ 2 Preliminaries
Multivariate Time Series Forecasting: Given a historical multivariate time series $X\in\mathbb{R}^{T\times N}$ of $T$ time points, MTSF aims to predict the next $F$ time points, i.e., $Y\in\mathbb{R}^{F\times N}$, where $F$ is called the forecasting horizon.
@@
- - - - - - - One-Fits-All Zhou et al. (2023) Foundation NIPS 2023 One-Fits-All
- - - - - - - $S^{2}$IP-LLM Pan et al. (2024) Foundation ICML 2024 $S^{2}$IP-LLM
- - - - - - - TEMPO Cao et al. (2024) Foundation ICLR 2024 TEMPO
@@
Transformer-based - - - - TQNet Lin et al. (2025) Specific ICML 2025 TQNet
@@
GNN-based - - - TimeFilter Hu et al. (2025) Specific ICML 2025 TimeFilter
@@
Asym.: Asymmetry, Lag.: Lagginess, Pol.: Polarity, Gw.: group-wise, Dyn.: Dynamism, and Ms.: Multi-scale. We will discuss them in Section 3.3.
@@ 3.1 Strategy Perspective
Channel Partiality (CP): The CP strategy strikes a balance between CI and CD, allowing each channel to retain a degree of independence while simultaneously interacting with other related channels. This approach emphasizes a hybrid state where channels selectively interact and exhibit partial correlations. Based on whether the number of related channels for each channel is fixed or dynamic, the existing CP methods can be divided into two categories: I) Fixed partial channels: These models fix the number of related channels for each channel, which means the set of associated channels remains constant over time. For example, in MTGNN Wu et al. (2020), the channel relationships are modeled as a K-regular graph, where each channel interacts with $K$ other channels using the CD strategy to model interdependencies, while the remaining channels interact through the CI strategy. Similarly, in MCformer Han et al. (2024b), each channel interacts with only $K$ other channels, maintaining the CI strategy with the rest to ensure computational efficiency and prevent overfitting. II) Dynamic partial channels: These models allow the number of related channels for each channel to be dynamic, changing over time and providing greater flexibility to adapt to varying scenarios. For instance, DUET Qiu et al. (2025) calculates channel similarity using metric learning in the frequency domain and then sparsifies the result. This creates a mask matrix, which is integrated into the attention mechanism of the fusion module, ensuring that each channel interacts only with relevant channels, reducing interference from noisy ones. Another example, CCM Chen et al. (2024), dynamically clusters channels based on their intrinsic similarities. To effectively capture the underlying time series patterns within these clusters, CCM utilizes a cluster-aware feed-forward mechanism, enabling tailored management and processing for each cluster.
@@ 3.2 Mechanism Perspective
Transformer-based: In recent years, Transformer has been widely applied to MTSF tasks, leveraging its powerful global modeling capability to effectively capture complex temporal dependencies and channel interactions. As shown in Figure 2, existing channel strategies based on the attention mechanism can be categorized into the following types. I) Naive Attention: These approaches all adopt the CD strategy, treating time series segments (patches) or the entire sequence of each channel as individual tokens, and directly applying attention mechanisms to model channel correlations. For instance, CARD Wang et al. (2024b) and iTransformer Liu et al. (2024b) represent the patches and series of each channel as independent tokens, respectively, and explicitly capture channel correlation using attention mechanisms. II) Router Attention: When the number of channels ($N$) is large, the computational complexity of channel attention reaches $O(N^{2})$, resulting in high computational costs. To address this, some methods propose optimization strategies to mitigate the computational complexity caused by the CD strategy. For example, Crossformer Zhang and Yan (2022) introduces a Router Mechanism for Naive Attention, which uses a small fixed number of $c$ “routers" ($c\ll N$) to gather information from all channels and redistribute it. This reduces the complexity to $O(2cN)=O(N)$. This mechanism effectively balances the modeling of channel correlation and computational efficiency. III) Frequency Attention: Some CD methods suggest that frequency-domain information is more effective for capturing inter-channel dependencies than time-domain information. For example, FECAM Jiang et al. (2023) transforms the time series data into the frequency domain and then employs Naive Attention to model inter-channel relationships in this domain. IV) Mask Attention: In naive attention, each channel calculates attention scores with all channels, which can be negatively affected by irrelevant channels. To mitigate this, Mask Attention provides an approach to avoid irrelevant noise by constructing a CP strategy. For example, DUET Qiu et al. (2025) generates mask matrices for Naive Attention, allowing each channel to focus on those beneficial for downstream prediction tasks, while mitigating the impact of noisy or irrelevant channels. This approach explicitly constrains the attention computation, improving the accuracy of channel correlation modeling.
@@
CNN-based: A Convolutional Neural Network (CNN) is a deep learning model that utilizes convolutional layers to extract local features from data. As illustrated in Figure 3, the explored CNN-based approaches use CNN Merging for channel modeling, such as Informer Zhou et al. (2021), Autoformer Wu et al. (2021), and FEDformer Zhou et al. (2022), use 1D convolution with a sliding operation along the temporal dimension in the initial feature extraction layers. These models treat different channels as distinct inputs to the convolution, whose features are then weighted and merged during the convolution process, enabling inter-channel interactions. Although TimesNet Wu et al. (2023a) employs 2D convolutions, it folds the temporal dimension into a 2D format, with channels still serving as independent input for weighted merging via convolution. Such models are all under the CD strategy.
GNN-based: By dividing the time series into different windows along time, where each channel within a window is treated as a node, and the correlations between channels are considered as edges, multivariate time series can be transformed into graph-based data. The GNN-based methods can be classified into dense and sparse graphs. In dense graphs, each node is typically connected to almost all other nodes, with the edges often representing the strength of correlation or the probability of correlated influence. Methods based on dense graphs, such as GTS Shang et al. (2021) and FourierGNN Yi et al. (2023), generally follow a CD strategy. In contrast, sparse graphs only retain necessary edges, with most nodes remaining independent. For instance, MTGNN Wu et al. (2020) preserves K edges per node, constructing a sparse K-regular graph. Different from this, MTSF-DG Zhao et al. (2023) sparsifies the adjacency matrix by filtering out low-probability edges based on a pre-set threshold. Methods based on sparse graphs belong to the CP strategy.
@@
The sparsity of the constructed graph determines whether the method follows a CD or CP strategy. Additionally, GNN-based models often rely on the type of graph they construct when implementing the CD or CP strategy. As shown in Figure 4, we classify graph types as follows: I) Simple Graph: A simple graph is the most basic graph model, where there is at most one edge between each pair of nodes. A well-defined graph structure is required for effective message passing. Researchers have used channel similarity metrics (MTGNN, MSGNet Cai et al. (2024), CrossGNN Huang et al. (2023a)) and data similarity metrics (GTS, WaveForM Yang et al. (2023)) to learn the correlation graph structure among multivariate channels. They utilize time-domain (MTGNN, MSGNet, CrossGNN, GTS) or frequency-domain (WaveForM) information as node learning features. Graph convolution-based message passing is applied within the simple graph to facilitate the transmission of dependency information among channels. II) Spatio-temporal Graph: Unlike a simple graph, a Spatio-temporal Graph incorporates multiple channels at different time steps into a single graph, further considering the relationships between channels across different time steps. This approach allows GNNs to simultaneously model both temporal and channel dependencies, effectively addressing potential compatibility issues between the temporal module and the GNNs. The main challenge of Spatio-temporal Graph-based methods is to address the efficiency issues in the graph construction and message passing stages. For example, FourierGNN uses fully connected graph construction and employs Fourier domain convolution operators to achieve a time complexity of $O(Nlog(N))$. Similarly, FC-STGNN Wang et al. (2024c) adopts the same graph construction method and employs moving-pooling convolution to achieve the same time complexity. III) Hyper Graph: Hypergraphs are an extension of graphs that allow hyperedges to connect multiple vertices, enabling the modeling of higher-order group interactions. Models based on hypergraphs assume that the interactions among channels are not pairwise but involve group-based interactions among multiple channels. Therefore, hypergraph-based models are inherently suitable for constructing CP strategies. ReMo Wu et al. (2023b) constructs multi-view and multi-scale hypergraphs, respectively, and designs message passing mechanisms on these hypergraphs to enable group-wise message propagation. It is noteworthy that they use different MLPs or clustering constraints to promote the expression of heterogeneity among groups. IV) Temporal Graph: In the real world, the correlation of time series data often changes over time, forming dynamic relational graphs. MTSF-DG and TPGNN Liu et al. (2022) use dynamic graphs and polynomial graphs, respectively, to model the variation patterns of these correlations. The CP model MTSF-DG combines historical and future relational graphs, leveraging memory networks and logical symbol learning to capture the impact of historical correlations on future correlations. The CD model TPGNN represents the correlation matrix as a matrix polynomial with time-varying coefficients to learn the evolving patterns of correlations.
@@
Others: In addition to the mechanisms mentioned above, some models have proposed alternative approaches. For example: I) The CD model SOFTS Han et al. (2024a) introduces the STAR module, which utilizes a centralized structure to first aggregate information from all channels using MLPs, and then distribute the aggregated information to each channel. This interaction not only reduces the complexity of inter-channel interactions but also minimizes reliance on individual channel quality. II) The CP model LIFT Zhao and Shen (2024) proposes a novel plugin, adaptable to all MTSF specific models, that efficiently estimates leading indicators and their lead steps at each time step. This approach enables lagging channels to utilize advanced information from a predefined set of leading indicators. III) C-LoRA Nie et al. (2024) introduces a channel-aware low-rank adaptation (C-LoRA) plugin, which is adaptable to all MTSF-specific models. It first parameterizes each channel with a low-rank factorized adapter to enable individualized treatment. The specialized channel adaptation is then conditioned on the series information to form an identity-aware embedding. Additionally, cross-channel relational dependencies are captured by integrating a globally shared CD model.
@@ 3.3 Characteristics Perspective
Lagginess: Lagginess refers to the fact that the current state of a certain channel is influenced by the past states of the other channels. Based on the lagginess characteristic, VCformer Yang et al. (2024) incorporates the joint effects of multi-step delays among channels when calculating the attention matrix. In contrast, FourierGNN Yi et al. (2023) and FC-STGNN Wang et al. (2024c) directly perform message passing between representations across different channels and time steps using spatiotemporal fully connected graphs. LIFT Zhao and Shen (2024), on the other hand, combines prior knowledge with neural network predictions to estimate the lag step.
@@
Group-wise: Group-wise refers to a phenomenon in which correlations among channels exhibit a grouping structure, characterized by strong correlations within the same group, weak correlations among different groups, and varying correlation dependencies across different groups. CCM Chen et al. (2024) and DUET Qiu et al. (2025) use clustering techniques to group channels for interaction, while ReMo Wu et al. (2023b) and Ada-MSHyper Shang et al. (2024) establish intra-group message passing through hyperedges. Furthermore, CCM and ReMo apply different MLPs for feature extraction within different groups. This approach facilitates the expression of differences among different groups.
Dynamism: The correlation among channels in multivariate time series is characterized by time-varying relationships, showing an overall dynamic change. First, methods based on MLP (such as TimeMixer Wang et al. (2024a), TTM Ekambaram et al. (2024)), where the weights remain constant across time steps, fail to capture dynamism. Methods that use a Transformer to consider channel correlation typically employ series tokens or patch tokens. Methods based on series tokens, such as iTransformer Liu et al. (2024b) and DUET Qiu et al. (2025), cannot capture dynamism. However, methods based on patch tokens, such as Crossformer Zhang and Yan (2022), assign different attention scores at different time patches, enabling the modeling of dynamism. In GNNs, only approaches where the graph structure changes over time can capture dynamism, such as MSGNet Cai et al. (2024), which computes the graph structure at each timestep. However, the aforementioned methods for modeling dynamism only consider different channel relationships at different time steps. In contrast, MSTF-DG Zhao et al. (2023), TPGNN Liu et al. (2022), and ESG Ye et al. (2022) propose that there is a direct connection between channel relationships across different time steps. For example, MSTF-DG uses previous channel relationships to directly infer the current channel relationships.
Multi-scale: Multi-scale refers to the phenomenon where the correlations among channels exhibit different behaviors at various time scales (such as hours, minutes, or seconds). MSGNet Cai et al. (2024) and Ada-MSHyper Shang et al. (2024) establish different graph structures across scales to describe the variations in correlation at different levels, and they achieve the fusion of correlation information at different scales through varying degrees of interaction. Considering the multi-scale heterogeneity of correlations helps the model better understand the multi-scale features of time series data, thereby generating more accurate predictions.
Figure 6: Method performance for varying correlation within datasets. Datasets are sorted by correlation (Weak → Strong).
Figure 7: Comparison of parameter counts and inference time for deep learning methods.